OpenStudy (anonymous):

simplify 3 divided by 5 minus 6i

OpenStudy (anonymous):

\[\frac{ 3 }{ 5-6i }\]=\[\frac{ 3(5+6i) }{ (5-6i)(5+6i) }\]=\[\frac{ 15+18i }{ 25-36i^2 }\]=\[\frac{ 15+18i }{ 25-(-36) }\]=\[\frac{ 15 }{ 61}\]+\[\frac{ 18 i}{ 61 }\]